(e^2x+1)+3-3(e^x+1)=e^x


1
Expert's answer
2021-06-08T04:21:02-0400

Solution.

"(e^{2x}+1)+3-3(e^x+1)=e^x; \newline
(e^{2x}+1) +3-3e^x-3=e^x; \newline
e^{2x}-4e^x+1=0; \newline"

Lets "y=e^x; \newline"

"y^2-4y+1=0; \newline
D=b^2-4ac; \newline
D=16-4*1*1=12; \newline
\sqrt{D}=\sqrt{12}=2\sqrt{3}; \newline
y_1=(-b+\sqrt{D})/2a; \newline
y_1=(4+2\sqrt{3})/2=2+\sqrt{3}=2+1.73=3.73; \newline
y_2=(-b-\sqrt{D})/2a; \newline
y_2=(4-2\sqrt{3})/2=2-\sqrt{3}=2-1.73=0.27; \newline"

The main natural logarithm formula is:

"ln(e^x)=x. \newline
y_1=e^{x_1};\newline
e^{x_1}=3.73;\newline
ln(e^{x_1})=ln(3.73);\newline
x_1*ln(e)=ln(3.73);\newline
x_1*1=ln(3.73);\newline
x_1=1.32; \newline
y_2=e^{x_2}; \newline
e^{x_2}=0.27; \newline
ln(e^{x_2})=ln(0.27); \newline
x_2=ln(0.27); \newline
x_2=-1.3."

Answer: "x_1=1.32; x_2=-1.3."